OpenStudy (anonymous):

What is the area of a sector with measure of arc equal to 90° and radius equal to 1 foot

OpenStudy (anonymous):

\[\pi/4\]

OpenStudy (anonymous):

because the radius is 1 foot, the area of the circle, with the formula \[\pi r^2\] = Area the area of the circle is \[\pi\] since the sector is 90degrees which is 1/4th, the area of the sector is pi/4

OpenStudy (anonymous):

so do that mean that it is 1/4 of 1 right which would be 0.25

OpenStudy (anonymous):

no, pi/4 because the whole circle is pi and you're only asked for 1/4th of the whole circle